kinetic energy of a particle moving alog a circle of radius R depends on distance covered as T=ks2 where k is constant. find the force acting on the particle a s function of s..

Grade:Select Grade

1 Answers

Sheetal
6 years ago
Given 1/2mv²=as²Differentiating both sides w.r.t. timeWe get. Tangential accn(w1) =2as/m Centripetal acceleration(w2) Now net accn=√[(w1)² + (w2)²] √[(2as/m)² + (2as²/mR)²] 2as/m √[1+(s/R)²]Now total force=m× net acceleration 2as√[(1+s²/R²)]